Sarah is two thirds as old as her cousin. The sum of their ages is 45. How old are Sarah and her cousin?
Answer by mananth(16946)

cousin = x years
Sarah = 2x/3 years
..
2x/3 +x = 45
LCD = 3
multiply by 3
2x/3 * 3 +3*x=45*3
2x+3x=135
5x=135
/5
x= 135/5
x=27 years cousins age.
..
Sarah = 2*27/3=18 years
..
m.ananth@hotmail.ca